Determine the point estimate of the population​ proportion, the margin of error for the following confidence​ interval, and the number of individuals in the sample with the specified​ characteristic, x, for the sample size provided.

Lower

boundequals=0.1290.129​,

upper

boundequals=0.5410.541​,

nequals=12001200

The point estimate of the population proportion is

nothing.

​(Round to the nearest thousandth as​ needed.)

The margin of error is

nothing.

​(Round to the nearest thousandth as​ needed.)

The number of individuals in the sample with the specified characteristic is

nothing.

​(Round to the nearest integer as​ needed.)
Answer #1

1)

The point estimate of the population proportion=(upper limit+lower limit)/2

=(0.541+0.129)/2

=0.335

2)

The margin of error is =(upper limit-lower limit)/2 ==(0.541-0.129)/2 =0.206

3)\number of individuals in the sample with the specified characteristic is=np

=1200*0.335

=402